# Flow-Aware Diffusion for Real-Time VR Restoration: Enhancing Spatiotemporal Coherence and Efficiency

### v1

@@ 3.7.2 Setting
Between different video viewing sessions, all participants were given a rest period of approximately one hour to ensure they had recovered from any symptoms of sickness caused by visual stimuli. Following each video viewing session, all participants were instructed to complete a standard Simulator Sickness Questionnaire (SSQ) [23].
@@ 3.7.3 Result
[…]
@@
[…]
@@ 4 Discussion
At the algorithmic level, the proposed U-MAD model integrates a motion-aware architectural design with a diffusion-based generative framework, demonstrating substantial improvements over conventional baselines such as U-Net Diffusion [50] and DIT [17]. This performance boost is primarily attributed to the incorporation of optical flow signals as explicit guidance during generation, which serve as strong priors for inter-frame motion estimation. Prior studies have emphasized the role of optical flow in maintaining motion continuity [27], and the results further validate this principle. In particular, U-MAD consistently improves temporal coherence and suppresses flicker and ghosting artifacts, especially in visually complex or fast-moving scenes. These improvements are evident in both quantitative metrics and user feedback. Participants in the user study reported noticeably reduced symptoms of dizziness, nausea, and disorientation, indicating that the motion consistency achieved by U-MAD is effective not just in algorithmic terms but also in enhancing real-world perceptual comfort. Compared to traditional approaches, such as field-of-view restriction [2], peripheral blurring [6], or scene simplification [14], the proposed method preserves visual fidelity, interaction freedom, and scene detail, thereby avoiding the trade-offs commonly associated with discomfort-mitigation techniques. Moreover, it generalizes effectively across procedurally generated and dynamic environments, reinforcing its utility in real-world deployment.
While the proposed approach demonstrates clear advantages, it also faces several practical challenges that warrant further discussion. The training process, while conducted offline, incurs significant computational cost due to the combination of diffusion-based generation and dense optical flow estimation. This resource requirement aligns with known challenges in deploying deep generative models for edge AI [11, 35]. To address this, we recommend a cloud-assisted pipeline wherein large-scale training is followed by compression techniques such as knowledge distillation, quantization, or structured pruning for efficient deployment on resource-constrained devices. Another practical consideration is inference latency, which is heavily influenced by the complexity of the optical flow estimator. While high-fidelity flow models enhance visual quality, they introduce substantial computational overhead, presenting a trade-off between performance and responsiveness. Additionally, optical flow alone may be insufficient for modeling complex scene dynamics such as object deformation, rapid camera motion, and occlusion.

### v2

@@ 3.7.2 Setting
Between different video viewing sessions, all participants were given a rest period of approximately one hour to ensure they had recovered from any symptoms of sickness caused by visual stimuli. Following each video viewing session, all participants were instructed to complete a standard Simulator Sickness Questionnaire (SSQ) [kennedy1993simulator].
@@ 3.7.3 Result
We conducted Paired Sample t-test and effect size for the metrics derived from the SSQ after verifying that the data satisfied assumptions of normal distribution (Shapiro-Wilk test, $p<.05$) and homogeneity of variance (Levene test, $p<.05$).
@@
The within-subject analysis revealed that the processed videos demonstrated significantly lower risk of inducing cybersickness compared to the original videos. As illustrated in Fig. 6, the analysis of the results reveals that the participants reported significantly lower ($t=3.7999,p<.01,Cohen^{\prime}s\text{ }d=.591$) overall SSQ socre after viewing the restored videos ($M=30.75,SD=25.61$) than viewing the original videos ($M=44.67,SD=21.29$). The participants reported significantly lower oculomotor ($t=5.1397,p<.01,Cohen^{\prime}s\text{ }d=.840$) after viewing the restored video ($M=24.00$, $SD=19.84$) than viewing the original videos ($M=39.58$, $SD=17.16$). The participants reported significantly lower ($t=2.4658,p=.0246,Cohen^{\prime}s\text{ }d=.471$) disorientation difficulty after viewing the restored videos ($M=40.99,SD=37.74$) than viewing the original videos ($M=57.23,SD=30.90$).
@@ 4 Discussion
At the algorithmic level, the proposed U-MAD model integrates a motion-aware architectural design with a diffusion-based generative framework, demonstrating substantial improvements over conventional baselines such as U-Net Diffusion [zhou2020unetredesigningskipconnections] and DIT [ho2022videodiffusionmodels]. This performance boost is primarily attributed to the incorporation of optical flow signals as explicit guidance during generation, which serve as strong priors for inter-frame motion estimation. Prior studies have emphasized the role of optical flow in maintaining motion continuity [koroglu2024onlyflowopticalflowbased], and the results further validate this principle. In particular, U-MAD consistently improves temporal coherence and suppresses flicker and ghosting artifacts, especially in visually complex or fast-moving scenes. These improvements are evident in both quantitative metrics and user feedback. Participants in the user study reported noticeably reduced symptoms of dizziness, nausea, and disorientation, indicating that the motion consistency achieved by U-MAD is effective not just in algorithmic terms but also in enhancing real-world perceptual comfort. Compared to traditional approaches, such as field-of-view restriction [alfano1990restricting], peripheral blurring [budhiraja2017rotation], or scene simplification [han2021deep], the proposed method preserves visual fidelity, interaction freedom, and scene detail, thereby avoiding the trade-offs commonly associated with discomfort-mitigation techniques. Moreover, it generalizes effectively across procedurally generated and dynamic environments, reinforcing its utility in real-world deployment.
While the proposed approach demonstrates clear advantages, it also faces several practical challenges that warrant further discussion. The training process, while conducted offline, incurs significant computational cost due to the combination of diffusion-based generation and dense optical flow estimation. This resource requirement aligns with known challenges in deploying deep generative models for edge AI [francy2024edgeaievaluationmodel, Loven2019EdgeAI]. To address this, we recommend a cloud-assisted pipeline wherein large-scale training is followed by compression techniques such as knowledge distillation, quantization, or structured pruning for efficient deployment on resource-constrained devices. Another practical consideration is inference latency, which is heavily influenced by the complexity of the optical flow estimator. While high-fidelity flow models enhance visual quality, they introduce substantial computational overhead, presenting a trade-off between performance and responsiveness. Additionally, optical flow alone may be insufficient for modeling complex scene dynamics such as object deformation, rapid camera motion, and occlusion.